THE METALLIC BOB OF A SIMPLE PENDULUM OF MASS 0.5 KG carries a positive charge of 0.25c. if the pendulum is suspended in an electric field of strength 15N C^-1DIRECTED from east to west, state the forces acting on the bob and derive the condition fo equilibrium. also find the angle subtended at the equilibrium. 9TAKE, g=10 m s^-2)

Grade:10

1 Answers

Saurav
5 years ago
FORCES ACTING : ELECTROSTATIC , WEIGHT & TENSION DUE TO STRING
ELECTROSTATIC FORCE=15 * 0.25 N FROM E TO W
WEIGHT=0.5 * 10 =5N
TENSION=T
LET ANGLE MADE WITH VERTICAL BE \Theta (CLOCKWISE FROM VERTICAL)
THE 
T sin(theta)=5N
T cos(theta)=15/4   N
therefore, tan(theta)=3/4
hence angle is 37 degress